OpenStudy (anonymous):

how to solve an equation using factoring and quadratic formula ??? 2x^ - 8x = 0 ???

OpenStudy (radar):

2x(x-4) x=0 x=4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

quadratic formula \[x = (-b + - \sqrt{b ^{2}-4ac)})/2a\]

OpenStudy (anonymous):

Using the formula. x= (\[-b \pm \sqrt{b) ^{2}-4ac}\])/2a a = 2 b= -8 c =0 . Plug in and Solve

OpenStudy (radar):

It factors easily enough.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks

OpenStudy (radar):

Both methods should result in identical solutions.

OpenStudy (anonymous):

oh okay

OpenStudy (radar):

That is a good check! Good luck.

OpenStudy (anonymous):

okay thank you